# ]https://www.c2n.universite-paris-saclay.fr/en/

### v1

@@ Abstract
Accurately determining the effective Hubbard parameter ($U_{eff}$) in Density Functional Theory plus U (DFT+U) remains a significant challenge, often relying on empirical methods or linear response theory, which frequently fail to predict accurate material properties. This study introduces BMach, an advanced Bayesian optimization algorithm that refines $U_{eff}$ by incorporating electronic properties, such as band gaps and eigenvalues, alongside structural properties like lattice parameters. Implemented within the Quantum Espresso platform, BMach demonstrates superior accuracy and reduced computational cost compared to traditional methods. The BMach-optimized $U_{eff}$ values yield electronic properties that align closely with experimental and high-level theoretical results, providing a robust framework for high-throughput materials discovery and detailed electronic property characterization across diverse material systems.
@@ II.1.1 Objective Function in BMach
Effectively utilizing Bayesian optimization hinges on the selection of a suitable objective function. In this instance, BMach’s surrogate objective function $f(\vec{U})$ encompasses not only the band gaps and eigenvalues but also the lattice parameters, as delineated in Eq. 4.
@@
In this optimized $U_{eff}$ value calculation, the ’Ref’ can be corresponded to the advanced hybrid functional-based HSE calculation or GW approximation-based G0W0 calculations. The weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are crucial as they dictate each term’s relative emphasis in the optimization process. These weights are meticulously chosen to reflect the relative importance of band gap accuracy, band structure fidelity, and lattice parameter congruence and are not necessarily constrained to sum up to unity (1); instead, they are selected based on the relative significance of each term. For example, if the bandgap error is considered more critical than the lattice parameter deviation, $\alpha_{1}$ might be larger than $\alpha_{3}$. The choice of these weights is informed by a combination of theoretical insights and empirical adjustments, ensuring that the algorithm adequately reflects the importance of each aspect of the material’s properties.
@@ II.2 Computational Tools and Techniques
InSb PAW, PBE+U (BMach-2D) -0.20 3.62 0.241 0.82 2.99 0.44
ONCV, PBE+U (BMach-2D) 3.80 2.90 0.239 0.82 2.84 0.46
ONCV, PBEsol+U (BMach-2D) 2.00 3.62 0.237 0.83 2.96 0.46
ONCV, PBEsol+U (BMach-1D) - 3.91 0.238 0.83 3.05 0.45
ONCV, PBEsol+U (LR) 0.85 2.98 0.070 0.82 2.87 0.45
ONCV, G0W0@HSE06 0.232 0.83 3.22 0.46
ONCV, HSEsol ($\mu=0.095$) 0.232 0.84 3.15 0.46
Experiment (temp. in K) [17] 0.235 (1.7) 0.81 (1.5) 3.02 0.42 (4)
@@
II.2.1 Ab initio DFT, HSE, G0W0, DFT+U Calculations
All ab initio structure relaxations and electronic band structure calculations were performed within the Density Functional Theory (DFT) framework using a Plane-Wave (PW) basis set and pseudopotentials-based method, as implemented in the Quantum Espresso (QE) toolkit [18]. Quasiparticle G0W0 calculations were carried out using the BerkeleyGW (BGW) package [19, 20].
Brillouin-zone integrations were executed using $\Gamma$-centred k-point meshes. For self-consistent field (SCF) calculations to determine the ground state charge density, and equilibrium lattice constants, an $8\times 8\times 8$ k-point grid was employed, yielding 29 irreducible k-points in the first Brillouin zone. Gaussian smearing with a smearing width of 0.005 Rydberg was used in SCF calculations. For non-self-consistent field (NSCF) calculations, a denser $16\times 16\times 16$ k-point grid was utilized. No symmetry operations were applied in all calculations to ensure the accuracy of band structure results. Furthermore, time-reversal symmetry ($k\rightarrow-k$) was not considered, thus treating the full $8\times 8\times 8$ Brillouin zone mesh as encompassing all 512 k-points as distinct, which desymmetrized the charge density and enhanced the precision of the computational results.
The band structures $E(k)$ were computed on a discrete k-mesh along high-symmetry directions, specifically for zincblende (ZB) InSb and InAs, from the Brillouin zone center $\Gamma$ (0,0,0) to the X point (0.5,0.0,0.5), W point (0.50,0.25,0.75), L point (0.5,0.5,0.5), back to $\Gamma$, and then to the K point (0.375,0.375,0.750) in crystal units. Fixed occupations were used to manage electron occupation in NSCF and band structure calculations.
Eigenvalues of the hybrid and G0W0 band structures were interpolated using the WANNIER90 code [21]. Considering the similar orbital character of the states near the band gap of narrow bandgap material InSb and InAs, the focus was on Wannierizing only the eight highest occupied and six lowest unoccupied bands around the band edges, employing sp3 initial projections. Spin-orbit coupling (SOC) corrections were applied to the interpolated bands for each eigenenergy $\varepsilon_{nk}$ and subsequently interpolated using maximally localized Wannier functions (MLWFs), following the method described by Malone and Cohen [22].
@@ III RESULTS AND DISCUSSIONS
((a)) 1D optimization results for InSb.
((b)) Band structure of InSb.
((c)) Band structure of InAs.
Figure 2: (a) shows the optimization results for the 1D model of InSb, including the GP mean prediction, observations, optimized point, and acquisition function after 100 BO steps. (b) and (c) show the band structures for InSb and InAs calculated using the DFT+U method with $U_{eff}^{Sb-5p}=3.91$ eV and $U_{eff}^{As-4p}=3.72$ eV, respectively. The $U_{eff}$ values are calculated using BMach with $\alpha_{3}=0$ to minimize the objective function with reference to the G0W0 band structures. Experimental lattice parameters of 6.479 Å for InSb and 6.058 Å for InAs are used.
@@
The BMach algorithm, as detailed in the methods section (see Figure 1), was employed to optimize the effective Hubbard $U$ parameters for narrow band gap materials such as Indium Antimonide (InSb) and Indium Arsenide (InAs).
III.1 Narrow band gap semiconductors
Narrow band gap materials are of significant interest due to their unique electronic properties, which make them suitable for various applications in optoelectronics and quantum computing. Among these materials, binary compounds such as Indium Antimonide (InSb) and Indium Arsenide (InAs) are notable for their small band gaps.
III.1.1 Indium Antimonide (InSb)
The effective Hubbard $U_{eff}$ parameters, band gaps ($E_{g}$), and spin-orbit splitting energies ($\Delta_{SO}$) for InSb using different computational setups are summarized in Table 1. The BMach-optimized $U_{eff}$ values yield band gaps and spin-orbit splitting energies that are in excellent agreement with experimental values and advanced theoretical methods.
The results of the linear response (LR) calculations, also presented in Table 1, highlight the limitations of the LR approach in accurately predicting the band parameters. Specifically, the $E_{g}$ values obtained using LR are significantly underestimated compared to experimental values and other advanced methods like BMach and G0W0@HSE06. For example, the LR calculation yields an $E_{g}$ of 0.070 eV for InSb, which is considerably lower than the experimental value of 0.235 eV. This discrepancy underscores the inadequacy of the LR method in this context.
Figure 2(a) illustrates the optimization results for the 1D model of InSb. The BMach algorithm efficiently converges to the optimal $U_{eff}$ values, providing accurate electronic and structural properties for InSb. The GP mean values offer a smooth approximation of the objective function, while the acquisition functions guide the sampling of new points, balancing exploration and exploitation to achieve accurate and reliable predictions.
Figure 2(b) shows the band structures of InSb calculated using G0W0, HSE, and DFT+U (1D and 2D) methods. The band structures obtained with BMach-optimized $U_{eff}$ values are in close agreement with those from G0W0 and HSE calculations, demonstrating the accuracy of the BMach approach.
III.1.2 Indium Arsenide (InAs)
Similar to InSb, the effective Hubbard $U$ parameters, band gaps ($E_{g}$), and spin-orbit splitting energies ($\Delta_{SO}$) for InAs are optimized using BMach. The results are summarized in Table 1, and the corresponding band structures are shown in Fig. 2(c).
@@ III.1.2 Indium Arsenide (InAs)
Figure 3 illustrates the optimization results for the 2D model of InAs. The BMach algorithm efficiently converges to the optimal $U_{eff}$ values, providing accurate electronic and structural properties for InAs. The GP mean values offer a smooth approximation of the objective function, while the acquisition functions guide the sampling of new points, balancing exploration and exploitation to achieve accurate and reliable predictions. The optimized 2D Hubbard $U$ values are -1.62 eV for In-5p and 3.86 eV for As-4p, as indicated by the asterisk in the figures.
Additionally, the margin of errors using BMach’s Bayesian Optimization (BO) is very low, and the prediction of energies such as $E_{g}$, $\Delta_{SO}$, $E_{g}^{\prime}$, and $\Delta_{SO}^{\prime}$ are very close to experimental and HSEsol or G0W0 results, further confirming the reliability and accuracy of BMach.
@@ III.2 Band Gap Optimization Results
The optimization results for the effective Hubbard $U$ parameters using the BMach algorithm are illustrated in Fig. 2(a) and Fig. 3. These figures show the Gaussian process (GP) mean values, standard deviations, and acquisition functions during optimization for the 1D and 2D models of InSb and InAs, respectively.
@@ IV CONCLUSIONS
The efficacy of BMach was demonstrated on narrow band gap semiconductors, specifically Indium Antimonide (InSb) and Indium Arsenide (InAs). The BMach-optimized $U_{eff}$ values yielded band gaps and spin-orbit splitting energies that align closely with experimental results and high-level theoretical methods such as G0W0 and HSE, confirming the algorithm’s reliability and accuracy. This alignment underscores the importance of considering both electronic and lattice parameters in the optimization process, providing a more comprehensive model for materials with strong electron correlation effects.
@@ Acknowledgements.
The computational work was conducted using the Ruche Supercomputing cluster at the Institute for Development and Resources in Scientific Computing (IDRIS), CNRS, Université Paris-Saclay. This research was supported by the ministère de l’enseignement supérieur, de la recherche et de l’innovation (MESRI) Doctoral grant. I would like to acknowledge Dr. Zhenglu Li (formerly at Lawrence Berkeley National Laboratory, University of California Berkeley, and currently at the University of Southern California) for the insightful discussion that inspired the direction of this research. Finally, I thank my colleagues at C2N Lab, CNRS, Université Paris-Saclay for their support.

### v2

@@ Abstract
Accurately determining the effective Hubbard parameter ($U_{eff}$) in Density Functional Theory plus U (DFT+U) remains a significant challenge, often relying on empirical methods or linear response theory, which frequently fail to predict accurate material properties. This study introduces BMach, an advanced Bayesian optimization algorithm that refines $U_{eff}$ by incorporating electronic properties, such as band gaps and eigenvalues, alongside structural properties like lattice parameters. Implemented for the Quantum ESPRESSO distribution, BMach demonstrates superior accuracy and reduced computational cost compared to traditional methods. The BMach-optimized $U_{eff}$ values yield electronic properties that align closely with experimental and high-level theoretical results, providing a robust framework for high-throughput materials discovery and detailed electronic property characterization across diverse material systems.
@@ II.1.1 Objective Function in BMach
Effectively utilizing Bayesian optimization hinges on the selection of a suitable objective function. In this instance, BMach’s surrogate objective function $f(\vec{U})$ encompasses not only the band gaps ($E_{g}$) and eigenvalues ($\epsilon$) but also the lattice parameters, as delineated in Eq. 4.
@@
In this optimized $U_{eff}$ value calculation, the $Ref$ can be corresponded to any advanced and accurate calculation such as HSE or GW calculations. The weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are crucial as they dictate each term’s relative emphasis in the optimization process. These weights are meticulously chosen to reflect the relative importance of band gap accuracy, band structure fidelity, and lattice parameter congruence and are not necessarily constrained to sum up to unity (1); instead, they are selected based on the relative significance of each term. For example, if the band gap error is considered more critical than the lattice parameter deviation, $\alpha_{1}$ might be larger than $\alpha_{3}$. The choice of these weights is informed by a combination of theoretical insights and empirical adjustments, ensuring that the algorithm adequately reflects the importance of each aspect of the material’s properties.
@@ II.2 Computational Tools and Techniques
II.2.1 Ab initio Calculations
All ab initio structure relaxations and electronic band structure calculations were performed within the Density Functional Theory (DFT) framework using a Plane-Wave (PW) basis set and pseudopotentials-based method, as implemented in the Quantum ESPRESSO (QE) distribution [18].
Brillouin-zone integrations were executed using $\Gamma$-centred k-point meshes. For self-consistent field (SCF) calculations to determine the ground state charge density, and equilibrium lattice constants, an $8\times 8\times 8$ k-point grid was employed. Gaussian smearing with a smearing width of 0.005 Rydberg was used in SCF calculations. For non-self-consistent field (NSCF) calculations, a denser $16\times 16\times 16$ k-point grid was utilized.
For a demonstrative use of BMach the band structures $E(k)$ for zincblende (ZB) indium arsenide (InAs) were computed on a discrete k-mesh along high-symmetry directions. Fixed occupations were used to manage electron occupation in NSCF and band structure calculations. Spin-orbit coupling (SOC) corrections were applied to the interpolated bands for each eigenenergy $\varepsilon_{nk}$ and subsequently interpolated using maximally localized Wannier functions (MLWFs), following the method described by Malone and Cohen [19].
@@ III RESULTS AND DISCUSSIONS
(a) 1D optimization results for InAs.
(b) Band structure of InAs.
Figure 2: (a) shows the optimization results for the 1D model of InAs, including the GP mean prediction, observations, optimized point, and acquisition function after 100 BO steps. (b) shows the band structure of InAs calculated using the DFT+U method with $U_{eff}^{As-4p}=3.72$ eV. The $U_{eff}$ values are calculated using BMach with $\alpha_{3}=0$ to minimize the objective function with the reference band structure. Experimental lattice parameter of 6.058 Å for InAs is used for all calculations.
@@
The BMach algorithm, as detailed in the methods section (see Section II.1 and Fig. 1), was employed to optimize the effective Hubbard $U$ parameters for narrow band gap material InAs.
III.1 Narrow band gap semiconductor
Narrow band gap materials are of significant interest due to their unique electronic properties, which make them suitable for various applications in optoelectronics and quantum computing. Among these materials, binary compound InAs is notable for its small band gap.
III.1.1 Indium Arsenide (InAs)
The effective Hubbard $U$ parameters for InAs are optimized using BMach. The band gaps ($E_{g}$) and spin-orbit splitting energies ($\Delta_{SO}$) are summarized in Table 1, and the corresponding band structures is shown in Fig. 2(b).
@@ III.1.1 Indium Arsenide (InAs)
Figure 2(a) illustrates the optimization results for the 1D model of InAs. The BMach algorithm efficiently converges to the optimal $U_{eff}$=3.72, providing accurate electronic and structural properties of InAs. The GP mean values offer a smooth approximation of the objective function, while the acquisition functions guide the sampling of new points, balancing exploration and exploitation to achieve accurate and reliable predictions.
Figure 3 illustrates the optimization results for the 2D model of InAs. The optimized 2D Hubbard $U$ values are -1.62 eV for In-5p and 3.86 eV for As-4p, as indicated by the asterisk ($\ast$) in the figures, providing accurate electronic and structural properties for InAs. The GP mean values offer a smooth approximation of the objective function, while the acquisition functions guide the sampling of new points, balancing exploration and exploitation to achieve accurate and reliable predictions.
Additionally, the margin of errors using BMach’s Bayesian Optimization (BO) is very low, and the prediction of energies such as $E_{g}$, $\Delta_{SO}$, $E_{g}^{\prime}$, and $\Delta_{SO}^{\prime}$ are very close to experimental results, further confirming the reliability and accuracy of BMach.
@@ III.2 Band Gap Optimization Results
The developement of the effective Hubbard $U$ parameters’ optimization process and its results using the BMach algorithm are illustrated in Fig. 2(a) and Fig. 3. These figures show the Gaussian process (GP) mean values, standard deviations, and acquisition functions during optimization for the 1D and 2D models of InSb and InAs, respectively.
@@ IV CONCLUSIONS
The efficacy of BMach was demonstrated on narrow band gap semiconductor indium arsenide (InAs) for this moment. Other materials are currently being tested under this framework, and the results will be updated in future versions of this manuscript. The BMach-optimized $U_{eff}$ values yielded band gaps and spin-orbit splitting energies that align closely with experimental results, confirming the algorithm’s reliability and accuracy. This alignment underscores the importance of considering both electronic and lattice parameters in the optimization process, providing a more comprehensive model for materials with strong electron correlation effects.
@@ Acknowledgements.
The computational work was conducted using the Ruche Supercomputing cluster at the Institute for Development and Resources in Scientific Computing (IDRIS), CNRS, Université Paris-Saclay. This research was supported by the French Ministry of Higher Education, Research and Innovationer (MESRI – France) doctoral grant. I would like to acknowledge Dr. Zhenglu Li (formerly at Lawrence Berkeley National Laboratory, University of California Berkeley, and presently at the University of Southern California) for the insightful discussion that inspired the direction of this research. Finally, I thank my colleagues at C2N Lab, CNRS, Université Paris-Saclay for their support.